your own portfolio that has $ 2650 invested in stock A and $4450 invested in stock B. If the expected return on these stocks are 8 percent and 11 percent, respectively, What is the expected return on the portfolio?
Answer #1

Total investment = 2650 + 4450

Total investment = 7,100

Expected return = (2650/7100)*0.08 + (4450/7100)*0.11

Expected return = 0.029859 + 0.068944

Expected return = 0.0988 or 9.88%